Use the square root property to solve:
2x^2 - 150 = 0

Answer by tutorcecilia(2152)

2x^2 - 150 = 0
2x^2-150+150=0+150 [solve for x]
2x^2=150
2x^2/2=150/2
x^2=75
sqrt(x^2)=sqrt(75) [take the square root of both sides]
x=8.66
.
Checking:
2(8.66)^2-150=0
150-150=0
0=0 [checks out]